Find the equations of the tangent and normal to each of the following conics, the lengths of the subtangent and subnormal, then trace the curve showing these lines.

y = x2 – 6x + 4 at (4, -4)


1
Expert's answer
2021-11-25T15:08:21-0500

tangent line:

"y-y_0=f'(x_0)(x-x_0)"

"f'(x)=2x-6"

"f'(4)=2"

"y+4=2(x-4)"

"y=2x-12"


normal line:

"y-y_0=-(x-x_0)\/f'(x_0)"

"y+4=-(x-4)\/2"

"2y+x=-4"


subtangent is the projection of the segment of the tangent onto the x-axis

subnormal is the projection of the segment of the normal onto the x-axis

A segment of a tangent to a curve lying between the tangency point (the point at which a tangent is drawn to a curve) and the intercept of the tangent with the x-axis is called the length of the tangent.

A segment of a line normal to a tangent lying between the tangency point and the intercept of the normal with the x-axis is called the length of the normal.



length of the subtangent: "6-4=2"

length of the subnormal: "4+4=8"